OpenStudy (anonymous):

Could you help me with a problem? r/3 + 6/3 = r/6 + 8/6 3r + 18 = r + 8 2r = -10 r = -5 The answer is incorrect. I need to know where I went astray.

OpenStudy (anonymous):

The second line

OpenStudy (anonymous):

I multiplied by six and got the wrong answer. What do I need to do in this next step?

OpenStudy (anonymous):

you hav to make the common denominator 6

OpenStudy (anonymous):

The common denominator would be six, correct?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

so: 2r/6 + 12/6 = r/6 + 8/6

OpenStudy (anonymous):

I was working on it when you beat me to it. Thank you very much for your help.

OpenStudy (anonymous):

Oh :D no problemo

OpenStudy (anonymous):

so did u get what r equals to?

OpenStudy (anonymous):

r=-4

OpenStudy (anonymous):

YUP :DD

OpenStudy (anonymous):

r/3 + 6/3 = r/6 + 8/6 Look to have easier equation if we want to have r should get out (/6) so multiplied all of the equation to6should have : 6*(r/3 + 6/3)=r/3 * 6/1 + 6/3 * 6/1 =2 r + 12 or 3r + 18 ?!:) And : (r/6 + 8/6) =r+8 So we have : 2 r + 12=r+8 and : 2r-r=+8-12=>r=-4 But finally answer there was (-5) Got it ?;) @gregatron I hope it can help you ;)